OpenStudy (anonymous):

find the slope -intercept equation for the line with indicated slope and y-intercept.... slope 6/7 , y-intercept (0,4)

OpenStudy (raffle_snaffle):

point slope formula

OpenStudy (anonymous):

eq of line in slope intercept form is given as y = mx + c we have slope m =6/7 , y-intercept c = y co-ordinate = 4 now plugging the values of m and c in the above eq we find \[y= mx+c \rightarrow y= \frac{6}{7} x+ 4 \rightarrow 7y = 6x+ 28\] i.e. 7y- 6x = 28 is the required eq in slope intercept form. @ken_boy05

OpenStudy (campbell_st):

the slope intercept form is y = mx + b your b value is b = 4 and m = 6/7 plug then in and you have the slope intercept form or the line

OpenStudy (anonymous):

y=mx+b formula

OpenStudy (anonymous):

i mean the slope -intercept equation

OpenStudy (anonymous):

so the correct answer will be identifying the values of m and b and substituting those values in the slope intercept equation

OpenStudy (campbell_st):

thats correct....